OpenStudy (anonymous):

A 75.0ml volume of 0.200M NH3 (kbase=1.8*10^-5) is titrated with 0.500M HNO3 . Calculate the pH after the addition of 21.0ml of HNO3.

OpenStudy (anonymous):

Apply Henderson's equation do you know it?

OpenStudy (anonymous):

The Henderson's equation is as shown below pH = pKa + log ([A-]/[HA]) [A-] = molar concentration of a conjugate base [HA] = molar concentration of a undissociated weak acid (M) The equation can be rewritten to solve for pOH: pOH = pKb + log ([HB+]/[ B ]) [HB+] = molar concentration of the conjugate base (M) [ B ] = molar concentration of a weak base (M) Example Problem Applying the Henderson-Hasselbalch Equation Calculate the pH of a buffer solution made from 0.20 M HC2H3O2 and 0.50 M C2H3O2- that has an acid dissociation constant for HC2H3O2 of 1.8 x 10-5. Solve this problem by plugging the values into the Henderson-Hasselbalch equation for a weak acid and its conjugate base. pH = pKa + log ([A-]/[HA]) pH = pKa + log ([C2H3O2-] / [HC2H3O2]) pH = -log (1.8 x 10-5) + log (0.50 M / 0.20 M) pH = -log (1.8 x 10-5) + log (2.5) pH = 4.7 + 0.40 pH = 5.1